Two charges are located on the positive ×-axis of a coordinate system. Charge q1 = 2.00x10-° C, and it is 0.020 m from the origin. Charge q2 =-3.00x10- C, and it is 0.040 m from the origin. What is the electric force exerted by these two charges on a third charge, q3 = 5.00x10-° C, located at the origin?


1
Expert's answer
2022-01-10T09:09:31-0500
"F=F_{13}+F_{23},\\\space\\
F=\dfrac1{4\pi \epsilon_0}\bigg(\dfrac{q_1q_3}{r_{13}^2}+\dfrac{q_2q_3}{r_{23}^2}\bigg)."


Please substitute the values with their signs according to the text of your problem because the values of charges are unclear.